OpenStudy (anonymous):

How do you illustrate the process of writing a function as a composition of simpler functions? Medal goes to the best answer.

OpenStudy (anonymous):

I have also searched in my book for examples of how to write a function such as w(x) = 1/t^2+2 as a composition of simpler functions. I am not finding any examples on this whole idea

OpenStudy (accessdenied):

We are mainly looking for simple parent functions to put together. Like: f(x) = 1/x g(x) = x^2 If we take f(g(x)), we put in g(x) for x in the definition of f(x). f(g(x)) = 1/g(x) = 1/x^2. Is that w(x) = 1/x^2 + 2, or w(x) = 1/(x^2 + 2) (its all in a denominator)?

OpenStudy (anonymous):

I am thrown off by the u(x)

OpenStudy (anonymous):

so my functions could be s(x) and t(x) = u(x)

OpenStudy (anonymous):

the normal flow I see is f(x) + g(x) = f(g(x))

OpenStudy (anonymous):

The original problem is write u(x) = 1/x^2-4 as a composition of simpler functions

OpenStudy (accessdenied):

Hm.. another question. I just need to know if this is: \( u(x) = \dfrac{1}{x^2} - 4 \) or \(u(x) = \dfrac{1}{x^2 - 4} \)

OpenStudy (anonymous):

Second one

OpenStudy (anonymous):

x^2-4 is in the denominator

OpenStudy (accessdenied):

Okay. So, note that we do have a parent of a reciprocal 1/x function. How can we tell? There is 1 on top and there is a function in the denominator. \( u(x) = \dfrac{1}{\color{blue}{x^2 - 4}} \) So what if we called one function g(x) = x^2 - 4, and another one f(x) = 1/x. Can you see how we might combine those to make u(x)?

OpenStudy (anonymous):

So you would have 1/x and x^2-4?

OpenStudy (accessdenied):

Yes, these are two simpler functions. I make up the function names f(x) = 1/x and g(x) = x^2 - 4 just so we can write it more conveniently. And we can compose the two functions in some way so that we get back u(x).

OpenStudy (anonymous):

I understand that part. Does it matter what we call the two function names?

OpenStudy (anonymous):

As long as we get back u(x)

OpenStudy (anonymous):

I was thinking there had to be some logical order like s(x) and t(x) gives you u(x)

OpenStudy (accessdenied):

It shouldn't. If you wanted to use that order, you could, but whether we call it f(x) or s(x) it will represent the same thing. :)

OpenStudy (anonymous):

This is starting to make sense. :)

OpenStudy (anonymous):

Now to find the domain because it is a square root, our domain has to be >= 0

OpenStudy (anonymous):

So you would set x^2 -4 as an inequality and solve for x?

OpenStudy (accessdenied):

Oh? it's a square root? Like, it was written... \( u(x) = \dfrac{1}{\sqrt{x^2 - 4}} \) ?

OpenStudy (anonymous):

Oh sorry... it was written as a fraction :)

OpenStudy (anonymous):

We set the bottom part of the fraction as an inequality... to find the domain?

OpenStudy (accessdenied):

The bottom part of a fraction can be positive or negative. However, it cannot strictly equal zero. So we want to solve an equality: Denominator = 0. These will be the points where denominator is 0, thus we would be dividing by zero (bad math).

OpenStudy (anonymous):

The way I wrote my answer is f(x) = 1/x, g(x) = x^2-4, u(x) = g(x^2-4) = 1/x / x^2 - 4

OpenStudy (accessdenied):

We would want to insert g(x) in for x for f(x), to get back to u(x). Like f(x) = 1/x f(3) = 1/3 f(x^2) = 1/(x^2) f(g(x)) = (f o g)(x) = 1/(g(x)) = 1/(x^2 - 4) [g(x) = x^2 - 4]

OpenStudy (anonymous):

when you have a composition of functions it just means that instead of the value x you insert a whole other function into the previous one e.g. if you have \[f(x)=x ^{2}\] and \[g(x)=2x\] f(g(x)) means that instead of the x in the function f you insert the function (g(x)) so f(g(x)) is really f(2x) which is equal to \[(2x)^{2}\]

OpenStudy (anonymous):

For the domain, it is going to be x^2 - 4 = 0

OpenStudy (accessdenied):

Well, the domain is going to be every real number except what you get by solving that equation!

OpenStudy (anonymous):

To solve that inequality, you would have x^2 - 4 = 0?

OpenStudy (anonymous):

No, because the denominator has to equal 0

OpenStudy (anonymous):

The domain of x^2-4 is all real numbers, (-oo, oo)

OpenStudy (anonymous):

Is that correct?

OpenStudy (accessdenied):

To clarify. The domain is actually all values \(x\) so that: \(x^2 - 4 \color{red}\ne 0 \). If \( x^2 - 4 = 0 \), then we have division by zero which is undefined. The domain of that function g(x) = x^2 - 4 is all real numbers, yes. The domain of u(x) = 1/(x^2 - 4) we do need to solve x^2 - 4 =/= 0 to remove values that do not work.

OpenStudy (anonymous):

You have x^2 not equal to 4?

OpenStudy (anonymous):

isn't the domain really all real numbers except for 2 and -2 because the expression equals 0 in those two cases

OpenStudy (accessdenied):

Yes, x^2 not equal to 4. Or x not equal to 2 or -2, when we take the square root (and considering both positive and negative roots).

OpenStudy (anonymous):

If you want to solve for x, then you take the square root of both sides...

OpenStudy (anonymous):

sqrt(x^2) - sqrt(4)

OpenStudy (anonymous):

x -2 not equal to 0... x is equal to 2

OpenStudy (anonymous):

Oh... 2 and -2

OpenStudy (accessdenied):

Yes. Taking a square root gives you both a positive and negative branch. Only because: \( (-2)^2 = 4 \text{ and } 2^2 = 4 \)

OpenStudy (anonymous):

Domain: (0, 2] U (-oo, -2)?

OpenStudy (anonymous):

I think I messed up somewhere

OpenStudy (accessdenied):

I think it would be easier to just say the domain is all real values of x where x not equal to 2 or -2. Interval notation, you need to have the interval from -oo to -2, from -2 to 2, and from 2 to oo.

OpenStudy (anonymous):

That is expressed as ||R, x not equal to 2, x not equal to -2

OpenStudy (anonymous):

I have used interval notation all the time that I don't know set builder that well

OpenStudy (accessdenied):

Oh, alright. Domain = { \( x \in \mathbb{R}\) | \(x \ne -2, 2\) } is one way. But if you are more comfortable with interval notation, we just split up the original interval at x=-2 and x=2. (-oo, oo) Breaks at x=-2... (-oo, -2) u (-2, oo) Also at x=+2 (-oo, -2) u (-2, 2) u (2, oo)

OpenStudy (anonymous):

Alright. That is what I have for interval notation, but I might start using set notation. I would give you 2 medals if I was able. You have given me a lot of help.

OpenStudy (accessdenied):

Medals and smart score are just superficial. I am always happy to help! :D

OpenStudy (anonymous):

Oh. That is great to hear! I want to be an ambassador and help out people too but I want to demonstrate a higher proficiency in math. :)

OpenStudy (accessdenied):

Both are nice goals to strive for. And with enough time and effort they are all possible. :) I know I spent a year or more studying Math on my own time before it just started clicking. It isn't always easy, but it is possible!

OpenStudy (accessdenied):

By Math, I speak mainly about Calculus I *

OpenStudy (anonymous):

I do appreciate the help. Thanks everyone

OpenStudy (anonymous):

might I just add that you ought to use square brackets before/after the + and -2 because that's how you denote that they're excluded from the domain whereas round brackets mean that they're included

OpenStudy (accessdenied):

Oh? I was pretty sure it was the other way around. Round brackets meant exclusion, and square brackets mean inclusion. [ a, b ] would mean "including endpoints a and b" to me

OpenStudy (anonymous):

I will chime in on this one. Brackets mean that it is excluded and ( mean included...

OpenStudy (anonymous):

Oh, actually @maida.beganovic is right

OpenStudy (anonymous):

[ ] a square bracket if you want to include the end value, or ( ) a round bracket if you don't

OpenStudy (anonymous):

That was your post. I think I got it mixed up. You had it right.

OpenStudy (accessdenied):

http://www.regentsprep.org/regents/math/algebra/ap1/intervalnot.htm I don't think I've ever seen it the other way with square means exclusion. But if you really wanted to be clear, you could just assert that () means exclusion/inclusion and [] the other. But you have to be consistent, above all.